# Fast nonparametric inference of network backbones for graph sparsification

### v1

@@ Abstract
A network backbone provides a useful sparse representation of a weighted network by keeping only its most important links, permitting a range of computational speedups and simplifying complex network visualizations. There are many possible criteria for a link to be considered important, and hence many methods have been developed for the task of network backboning for graph sparsification. These methods can be classified as global or local in nature depending on whether they evaluate the importance of an edge in the context of the whole network or an individual node neighborhood. A key limitation of existing network backboning methods is that they either artificially restrict the topology of the backbone to take a specific form (e.g. a tree) or they require the specification of a free parameter (e.g. a significance level) that determines the number of edges to keep in the backbone. Here we develop a completely nonparametric framework for inferring the backbone of a weighted network that overcomes these limitations by automatically selecting the optimal number of edges to retain in the backbone using the Minimum Description Length (MDL) principle from information theory. We develop two encoding schemes that serve as objective functions for global and local network backbones, as well as efficient optimization algorithms to identify the optimal backbones according to these objectives with runtime complexity log-linear in the number of edges. We show that the proposed framework is generalizable to any discrete weight distribution on the edges using a maximum a posteriori (MAP) estimation procedure with an asymptotically equivalent Bayesian generative model of the backbone. We compare the proposed method with existing methods in a range of tasks on real and synthetic networks.
@@ I Introduction
In a range of practical applications including the simulation of epidemics or information cascades mercier2022effective ; mathioudakis2011sparsification and the visualization of large networks imre2020spectrum , it is often helpful to alleviate computational burden by reducing the number of edges in a network while preserving key properties of interest for computations. This task is known as graph sparsification spielman2008graph , and often involves identifying a network backbone consisting of a network subgraph in which important characteristics such as degree and strength distributions serrano2009extracting or shortest paths grady2012robust are preserved. As there are many possible objectives for sparsification, there have been many proposed methods for identifying backbones in networks neal2022backbone ; yassin2023evaluation . Some of these methods can be used to sparsify unweighted graphs, but they most often applied to weighted graphs in which edges with higher weight are more likely to be kept in the network backbone when the weight signals edge importance.
@@
A number of existing methods for graph sparsification are stochastic, aiming to preserve properties such as network spectra with high probability when sampling edges based on their structural properties spielman2011spectral ; su2024generic ; john2017single . However, most network backboning methods are deterministic in nature and prune the network down to a single subgraph for analyses. Network backboning methods can generally be categorized as either global or local in nature yassin2023evaluation . Global backboning methods consider the whole structure of the network when considering whether or not to remove an edge, while local backboning methods only consider the neighborhood surrounding the nodes at one or both ends of the edge in question.
@@
The simplest method for network backboning is a global thresholding procedure, in which all edges with weight less than some pre-defined threshold are dropped from the network. However, this has been shown to poorly preserve many structural properties of interest, especially surrounding nodes with generally weak ties serrano2009extracting . Another straightforward global backboning method is to compute the minimum (maximum) spanning tree with respect to weights transformed to represent some notion of distance (similarity) between nodes wu2006transport . Spanning trees have many useful applications but are naturally acyclic so cannot possibly preserve important network properties such as transitivity, clustering, or path lengths. The High Salience Skeleton (HSS) introduced in grady2012robust is a global backboning method that aims to identify edges critical for routing along shortest paths in the network. This method has the appealing properties of being principled and relatively insensitive to its only user-chosen threshold, but has a computational burden that is at least quadratic in the number of nodes in practice and often does not preserve the global connectivity of the graph (as we will see in Sec. III.2). Other global backboning methods tend to suffer from the same issue of scalability, as they often require the computation of shortest paths zhang2018extracting ; simas2021distance or community structure as input rajeh2022modularity .
@@
Local backboning methods provide a scalable alternative to global methods by considering edges based only on nearby network structure, with the tradeoff that they may fail to preserve large-scale properties. One of the most popular methods for local network backboning is the Disparity Filter serrano2009extracting , which drops edges if their weight falls below a pre-specified significance level under a suitable null model relative to the other edges in a node’s neighborhood. The Polya Filter marcaccioli2019polya generalizes the disparity filter to a whole family of Polya urn null models, allowing for greater flexibility in modeling weight heterogeneity. A number of other statistical methods have also been considered that aim to remove edges that are deemed insignificant under a suitable null model tumminello2011statistically ; dianati2016unwinding ; casiraghi2017relational which may be distribution-free foti2011nonparametric . All of these methods have the requirement of a user-defined threshold or significance level that directly determines the number of edges present in the final backbone. While this approach can be advantageous for identifying multiple scales at which a network exhibits significant structural patterns serrano2009extracting , it also leaves arguably the most critical backboning choice up to the user and does not have a natural mechanism for selecting the single backbone representation to be used in an application.
@@
Properly regularized Bayesian methods for network reconstruction have circumvented the need for fixing free parameters when inferring statistically significant network structure newman2018network ; peel2022statistical ; peixoto2024network . These methods are statistically principled and robust, but are usually not suitable for backboning applications because they often increase the number of edges in a network when there is strong evidence for edges being omitted during the measurement process. They also tend to be computationally expensive due to the need to search over edge configurations involving node pairs that do not have any observed edge, the number of which scales quadratically with the network size. There are fast approximations aimed to alleviate this issue and provide better asymptotic runtime scaling, peixoto2024scalable , although the resulting scaling is still worse than that of many backboning methods due to the inherently higher complexity of the reconstruction problem.
@@
Here we develop a fast, completely parameter-free method for inferring the backbone of a weighted network based on the Minimum Description Length (MDL) principle from information theory rissanen1978modeling . Our description length objective computes the amount of information required to transmit the edge weights for a given network topology using a high weight backbone as an intermediate step. Minimizing over this encoding with an asymptotically optimal greedy procedure provides the backbone that best compresses the edge weight structure in the network, while automatically selecting for the number of edges to retain in the backbone. We develop two encodings using our framework—a global encoding and a local encoding—among which we can rigorously select an optimal backbone by comparing the final description length values. We show that the MDL framework is generalizable to any discrete edge weight distribution by reframing the problem as a Bayesian maximum a posteriori estimation procedure, allowing for a generative interpretation of the backbone formulation. Our method provides a flexible, principled, interpretable solution for network backboning that easily scales to large networks. We test our method on a range of real and synthetic network data, finding that it compares favorably against existing methods in a number of dimensions while eliminating the need for tuning any parameters.
@@ II Methods
The method we derive here for network backboning is motivated by the Minimum Description Length (MDL) principle, which states that the optimal model for a dataset is the model that allows us to transmit the dataset using the fewest bits rissanen1978modeling . By considering both the information to transmit a dataset given a model and the information to transmit the model itself, the MDL principle provides a flexible framework for penalizing overly complex models to reduce the risk of overfitting. The MDL principle has been employed in the network context for a number of methods in the past, for the tasks of community detection peixoto2019bayesian ; kirkley2022spatial ; peixoto2023implicit , graph comparison coupette2021graph ; coupette2022differentially , graph summarization liu2018graph ; kirkley2023compressing , and for extracting structural regularities in dynamic networks and hypergraphs shah2015timecrunch ; young2021hypergraph ; kirkley2024identifying ; kirkley2024inference . In the backboning context, the MDL principle allows us to naturally promote sparsity in the inferred backbone while encouraging the backbone to carry the most significant weights in the network.
We consider two encodings for network weights that utilize a backbone to compress the weight information in the edges of the network. We first develop a global encoding and then apply similar reasoning to adapt this encoding to transmit node neighborhoods individually. We demonstrate that both encodings can be minimized using fast greedy procedures, allowing us to easily identify the MDL-optimal backbone configurations. We then demonstrate the equivalence of our MDL formulation with MAP estimation using a Bayesian generative model for the edge weights, allowing for generalization of our framework to a broader class of weight distributions on the edges.
II.1 Global Network Backboning Algorithm

### v3

# Fast nonparametric inference of network backbones for weighted graph sparsification

@@ Abstract
Network backbones provide useful sparse representations of weighted networks by keeping only their most important links, permitting a range of computational speedups and simplifying network visualizations. A key limitation of existing network backboning methods is that they either require the specification of a free parameter (e.g. significance level) that determines the number of edges to keep in the backbone, or impose specific restrictions on the topology of the backbone (e.g. that it is a spanning tree). Here we develop a completely nonparametric framework for inferring the backbone of a weighted network that overcomes these limitations and automatically selects the optimal set of edges to retain using the Minimum Description Length (MDL) principle. We develop objective functions for global and local network backboning which evaluate the importance of an edge in the context of the whole network and individual node neighborhoods respectively and are generalizable to any weight distribution under Bayesian model specifications that fix the average edge weight either exactly or in expectation. We then construct an efficient and provably optimal greedy algorithm to identify the backbone minimizing our objectives, whose runtime complexity is log-linear in the number of edges. We demonstrate our methods by comparing them with existing methods in a range of tasks on real and synthetic networks, finding that both the global and local backboning methods can preserve network connectivity, weight heterogeneity, and spreading dynamics while removing a substantial fraction of edges.
@@ I Introduction
In a range of practical applications including the simulation of epidemics or information cascades mercier2022effective ; mathioudakis2011sparsification and the visualization of large networks imre2020spectrum , it is helpful to alleviate computational burden by reducing the number of edges in a network while preserving key properties of interest for computations. This task is known as graph sparsification spielman2008graph , and often involves identifying a network backbone consisting of a network subgraph in which important characteristics such as degree and strength distributions serrano2009extracting or shortest paths grady2012robust are preserved. As there are many possible objectives for sparsification, there have been many proposed methods for identifying backbones in networks neal2022backbone ; yassin2023evaluation . Some of these methods can be used to sparsify unweighted graphs, but they most often applied to weighted graphs in which edges with higher weight are more likely to be kept in the network backbone when the weight signals edge importance with respect to the network structure or dynamics.
@@
A number of existing methods for graph sparsification are stochastic, aiming to preserve properties such as network spectra with high probability when sampling edges based on their structural properties spielman2011spectral ; su2024generic ; john2017single . However, most network backboning methods are deterministic in nature and prune the network down to a single subgraph for analyses. Network backboning methods can generally be categorized as either global or local in nature yassin2023evaluation . Global backboning methods consider the whole structure of the network when considering whether or not to remove an edge, while local backboning methods only consider the neighborhood surrounding the node(s) at one or both ends of the edge in question.
@@
The simplest method for weighted network backboning is a global thresholding procedure in which all edges with weight less than some pre-defined threshold are dropped from the network. Another straightforward global backboning method is to compute the minimum (maximum) spanning tree with respect to weights transformed to represent some notion of distance (similarity) between nodes wu2006transport . Spanning trees have many useful applications but are naturally acyclic so cannot possibly preserve important network properties such as transitivity, clustering, or path lengths. The High Salience Skeleton (HSS) introduced in grady2012robust is a global backboning method that aims to identify edges critical for routing along shortest paths in the network. This method has the appealing properties of being principled and relatively insensitive to its only user-chosen threshold, but has a computational burden that is at least quadratic in the number of nodes in practice and often does not preserve the global connectivity of the graph (as we will see in Sec. III.2). Percolation thresholding, which imposes the maximum global weight threshold that allows the network to be a single connected component li2007transport ; haimovici2013brain , provides an alternative global backboning method which alleviates this issue at the cost of often requiring a large fraction of edges to be retained in order to ensure all nodes are connected. Other global backboning methods tend to suffer from the same issues of scalability and/or placing tight restrictions on the backbone topology, with scalability a particular issue for methods that require the computation of shortest paths zhang2018extracting ; simas2021distance or community structure rajeh2022modularity .
@@
Local backboning methods provide a scalable alternative to global methods by considering edges based only on nearby network structure, with the tradeoff that they may fail to preserve large-scale properties. One of the most popular methods for local network backboning is the Disparity Filter serrano2009extracting , which drops edges if their weight falls below a pre-specified significance level under a suitable null model relative to the other edges in a node’s neighborhood. The Polya Filter marcaccioli2019polya generalizes the disparity filter to a whole family of Polya urn null models, allowing for greater flexibility in modeling weight heterogeneity. A number of other statistical methods have also been considered that aim to remove edges that are deemed insignificant under a suitable null model tumminello2011statistically ; dianati2016unwinding ; casiraghi2017relational ; foti2011nonparametric . All of these methods have the requirement of a user-defined threshold or significance level that directly determines the number of edges present in the final backbone. While this approach can be advantageous for identifying multiple scales at which a network exhibits significant structural patterns serrano2009extracting , it also leaves arguably the most critical backboning choice (the number of edges to retain) up to the user and does not have a natural mechanism for selecting the single backbone representation to be used in an application.
@@
Properly regularized Bayesian methods for network reconstruction have circumvented the need for fixing free parameters when inferring statistically significant network structure newman2018network ; peel2022statistical ; peixoto2025network . These methods are statistically principled and robust, and are in part aimed at addressing concerns that backboning and thresholding methods—by enforcing the removal of edges—may impose unwanted biases on the structure of the pruned networks and remove important connectivity information cantwell2020thresholding . Despite their range of benefits, network reconstruction methods are usually not suitable for network backboning because they often increase the number of edges in a network when there is strong evidence for edges being omitted during the measurement process. They also tend to be computationally expensive due to the need to search over edge configurations involving node pairs that do not have any observed edge, the number of which scales quadratically with the network size. There are fast approximations aimed to alleviate this issue and provide better asymptotic runtime scaling peixoto2024scalable , although the resulting scaling is still worse than that of many backboning methods due to the inherently higher complexity of the reconstruction problem. Network backboning methods’ comparatively low complexity and prioritization of sparsity in the final network representation are often advantageous compared to network reconstruction techniques when considering visualization and downstream tasks with high computational overhead yassin2023evaluation .
There are also important classes of methods aimed at sparsifying networks so as to preserve or destroy one particular structural or dynamical property. For example, spectral sparsification spielman2008graph aims to sparsify a network while preserving the spectral properties of a particular graph operator such as the graph laplacian or graph adjacency matrix. And network dismantling braunstein2016network ; zdeborova2016fast aims to identify the minimal set of nodes or edges to remove in order to destroy any extensive connected components. While these methods are more effective than general network backbones for graph sparsification in specific problem contexts, network backbones are useful representations when a variety of exploratory and simulation-based analyses are performed on the same network, and the goal is to preserve connectivity, weight heterogeneity, and other structural and dynamical properties while reducing the dataset size for simpler manipulation and computation. As such, evaluation of network backbones is intrinsically multifaceted yassin2023evaluation .
@@
Here we develop fast, completely parameter-free methods for inferring global and local backbones of a weighted network based on the Minimum Description Length (MDL) principle rissanen1978modeling . We adapt our framework to Bayesian model specifications in which total edge weight is fixed in expectation (“canonical” models, informally) and exactly (a “microcanonical” model), proving that simple greedy approaches can find the exact optimum in a subset of these models. Minimizing over these objectives provides the backbone that best compresses the edge weight structure in the network, while automatically selecting for the number of edges to retain in the backbone. Our methods provide flexible, principled, and interpretable solutions for network backboning that easily scale to large networks. We test our methods on a range of real and synthetic network data, finding that they compare favorably against existing methods in a number of dimensions while eliminating the need for tuning any parameters. Finally, we study an example application in network epidemic spreading, finding that our methods allow for the accurate estimation of the associated percolation transition on networks while substantially reducing the required runtime for the calculation.
@@ II Methods
We start by deriving a general global backboning objective through maximum a posteriori (MAP) estimation in a Bayesian generative model of the network weights in which the total network weight is only fixed in expectation. We informally call this model a canonical model to emphasize that the exact total weight is not known ahead of time. (A truly “canonical” model is more restrictive as it requires a maximum entropy distribution with weight fixed in expectation, which in the present case is only true for the geometric and exponential distributions for integer-valued and real-valued weights respectively.) We apply this generative model to the level of node neighborhoods to develop a local backboning model.
We then discuss the approximate equivalence between our canonical Bayesian generative model in the case of geometric weights and a simpler microcanonical Bayesian formulation that is fixes the total network weight to an observed value. (Here we only focus on a truly “microcanonical” model that is uniform over all positive integer edge weight configurations with fixed total sums in the backbone and non-backbone.) Finally, we demonstrate that our microcanonical objective permits a fast, exact greedy procedure for identifying the MDL-optimal backbones, and extend our proof of greedy optimality to a set of common weight distributions under the canonical model formulation.
@@
Our methods are motivated by the Minimum Description Length (MDL) principle, which states that the optimal model for a dataset is the model that allows us to transmit the dataset using the fewest bits rissanen1978modeling . By considering both the information to transmit a dataset given a model and the information to transmit the model itself, the MDL principle provides a flexible framework for penalizing overly complex models to reduce the risk of overfitting. The MDL principle has been employed in the network context for a number of methods in the past, for the tasks of community detection peixoto2019bayesian ; kirkley2022spatial ; peixoto2023implicit , graph comparison coupette2021graph ; coupette2022differentially , graph summarization liu2018graph ; kirkley2023compressing , and for extracting structural regularities in dynamic networks and hypergraphs shah2015timecrunch ; young2021hypergraph ; kirkley2024identifying ; kirkley2024inference . In the backboning context, the MDL principle allows us to naturally promote sparsity in the inferred backbone while encouraging the backbone to carry the most significant weights in the network. There are many possible weight distributions for our canonical backboning models, each resulting in a different MDL formulation, and so technically we are considering not one but many possible MDL formulations of the backboning problem.
II.1 Canonical Models for Global and Local Backboning